ALFA (lband multi-beam system).

june,2002

     The Arecibo Lband Feed Array (alfa) consists of 7 feeds. 6 hexagonal feeds with 1 in the center.


Sections


History
recent system performance measurements
Calibration
measurements

Tsys
Monitoring alfa performance with users data Alfa horn locations on the rotary floor Jumps in totalPwr from Radars
Measurements in date order
IF LO system
alfa,pdev schematic
bandpass ripples from fiber

pre installation motor controller
alfa shutter
cal values rfi measurements


History


Calibration measurements (top)  (see also sysperf alfa)

30may16: calibration scans to check alfa gain,pointing after installation.
16may15: calibration scans to check alfa gain,pointing after installation
27may14: calibration scans after alfa reinstalled
05feb14: alfa bandpass (uncalibrated and in deg K)
03may12: calibration scans after alfa reinstalled 20apr12.
sep_oct11: calibration scans after alfa reinstalled.
02nov09: calibration scans after alfa re installed, alfa rotation angle offset check.
29jun09: calOn,calOf: averaging over the band vs linear fits.
09may05:comparing the system performance of alfa and lbw at OH frequencies (with10db amps in)
01apr05:comparing the system performance of alfa and lbw at OH frequencies (no 10db amps in)
16jan05: beammaps on 3C138 using calibration scan data.
29nov04: calibration scans on B1756+134. Beammaps and Measurement of the zero offset in the rotation angle.
16aug04: system performance (spider scans) at 1665 Mhz
29jul04: measuring the cal values at 1665 Mhz using sky and absorber.
22apr04: alfaTsys/Tcal using the old if/lo system.
apr04: Alfa cal values measured in the lab.


Tsys  (top)


Monitoring alfa performance


Jumps in Tsys caused by radars.  (top)

Note: the 26aug08 measurements show that the long recovery time caused when the radars hit alfa are a gain change in the dewar..

10oct12: alfalfa data sep12 shows long recovery time in beam1b. (new amps installed apr12)
19nov09: alfa recovery with the leds installed
24sep08  alfa long recovery times vs alfa rotation angle
23sep08 alfa long recovery times after bias changes (on sky).
18sep08:remove zener bias circuit 3a amp3, modify Vd,I for 3a and 3b.
16sep08:remove zener bias circuit 3b amp3, modify Vd,I.
11sep08: pulsed noise into alfa horn, modifying the bias currents.
26aug08:Pulsed noise into alfa horn shows long recovery is in dewar
04apr08: pulsed noise source into alfa mixer to create long recovery times.
25mar08: za location for power jumps along the meridian (using pdev)
25mar08: faa radar causes jumps and slow recovery. measured with pdev spectrometer.
A2010 total power strips show jumps: see 11jan07,13jan07,08mar06, 28apr06.. etc
06mar08: radar compression causes jumps and slow recovery. measured with pdev spectrometer.
01aug05: compression, baseline variation with long recovery time in p2030 data (from faa radar)
28feb,03mar05: faa radar causes increase in power with slow recovery time.
28sep04: compression of alfa receiver from faa radar. Az dependence (64Usecsampling)
24sep04: radar causes power jumps in some alfa beams.
06sep04: jumps in the total power while doing drift scans (A1946 precursor).  

Bandpass ripples caused by the fiber.

18oct11: 174khz ripple in bandpass from alfa fiber (data taken 14oct11)
23apr11: ripples in bandpass caused by fiber.
01jun05: the 174 khz ripple in pixel 4A of alfa.
21may05: pixel 4a still has a173Khz ripple (amplitude of .09K).
08sep04: 174 khz ripple in pixel4a alfa data.


Measurements by date  (top)

30may16: calibration scans to check alfa gain,pointing after installation.
22jan16: bm6b instabilities
16may15: calibration scans to check alfa gain,pointing after installation
27may14: calibration scans after alfa reinstalled
05feb14: alfa bandpass (uncalibrated and in deg K)
15jan14: alfa winking cal show total power time variation with calon, caloff.
10oct12: alfalfa total power data shows long recovery times for beam 1b sep12
14jun12: 20 hz birdies from alfa monitoring come from the bias monitoring.
12jun12: p2030 reported 1.5 hz rfi in data (not present). .838 hz rfi is visible.
02jun12: glitches in alfa gain caused by alfa dewar monitoring,.
03may12: calibration scans after alfa reinstalled 20apr12.
07dec11: run switching test on alfa filter bank.
28nov11: beam2a shows some jumps in ages data.
27oct11: fit C0,1az,3az terms to alfa pointing errors.
sep_oct11: calibration scans after alfa reinstalled.
19oct11: cleaning alfa fibers caused beam1A tsys to jump to 150K.
18oct11: 174khz ripple in bandpass from alfa fiber (data taken 14oct11)
10oct11: tsys beam4a while dana uses the spectrum analyzer.
07oct11:a2010 total power for 06-07oct11 showing bm4awith low drainI and bm6b failures.
03oct11: beam6b power drops out , becomes erratic.
27sep11: beam5a dropout seen by alfalfa.
26jun11: a2130 adc,pfb overflows vs za for jun11.
20may11:a2130 caldeflection beam 1b jumps occurs with 20k diode temp jump.
04may11: a2130(galfacts) bm4B caldeflection has za dependence
02may11: a2130 (galfacts) cal deflection show gain stability.
26apr11: instabilities  in beam 1b.
23apr11: ripples in bandpass caused by fiber.
28mar11: radar intermods in galfa spectrometer.
16mar11: plots of alfa temps during cooldown sequence (.ps) (.pdf)
08mar11- alfa motor controller, dome cameras rfi at 750 Mhz.
02mar11: 327 rfi caused by alfa motor controller.
15feb11: 5.2 sec oscillation in beam 2 from the dewar bias monitoring.
30nov09: Alfa outer beam position errors in mock fits files.
19nov09: 2.5 Mhz oscillations from bias box
19nov09: look at total power stability with leds on.
19nov09: alfa recovery with the leds installed
02nov09: calibration scans beam 0 after alfa re installed.
11sep09: 1390 rdr appears in alfalfa data.
29jun09: calOn,calOf: averaging over the band vs linear fits.
14may09: Intermods seen in alfa/pdev
14may09: table of intermods created in 1st and 2nd mixers of alfa/pdev.
26apr09: radar looking rfi in alfa data at 1403,1424, 1475 Mhz.
11apr09: gps L3 in a2010 data jan09-mar09.
19mar09: GPS L2, GLONASS L2, and COMPASS E6 rfi 1200-1300 MHz
24sep08  alfa long recovery times vs alfa rotation angle
23sep08 alfa long recovery times after bias changes (on sky).
21sep08:check alfa stability after bias voltages changed.
18sep08:remove zener bias circuit 3a amp3, modify Vd,I for 3a and 3b.
16sep08:remove zener bias circuit 3b amp3, modify Vd,I.
11sep08: pulsed noise into alfa horn, modifying the bias currents.
04sep08: gps satellite flies thru a2010 observation causing birdies at 1346.7
26aug08:Pulsed noise into alfa horn shows long recovery is in dewar
04apr08: pulsed noise source into alfa mixer to create long recovery times.
25mar08: za location for power jumps along the meridian (using pdev)
25mar08: faa radar causes jumps and slow recovery. measured with pdev spectrometer.
09may05:comparing the system performance of alfa and lbw at OH frequencies (with10db amps in)
03apr07: radar comb seen in 28mar07 data using serendip 5 and alfa.
12nov06: alfa beam1a is noisier than the other beams. Probaby coming from the fiber transmitter.
18apr06: one of alfa's refrigerators fails.
01apr05:comparing the system performance of alfa and lbw at OH frequencies (no 10db amps in)
24feb06: Harmonics created in the alfa/wapp configuration with cfr=1385Mhz.
21feb06: measure the channel number of the center of the wapp/alfa spectra.
17oct05:   location of 1350 faa radar aliases in alfa.
15sep05:  Measurements show that the cal deflection is not a function of the rotation angle.
27aug05: the alfa motor controller was creating the 16Khz comb in the 327 Mhz rcvr.
22aug05: looking at TsysA-TsysB vs azimuth and feed rotation angle.
01aug05: compression, baseline variation with long recovery time in p2030 data (from faa radar)
19jul05: galfa pixel4a,6a were switched (11jul05 to 19jul05).
13jul05: power levels in galfa pixel 2b changing.
01jun05: the 174 khz ripple in pixel 4A of alfa.
12jun05: 1428 Mhz biridie in alfa 18may05 thru 23may05
31may05: galfa pixel4a,6a were switched (before 02may05 until 31may05).
21may05: pixel 4a still has a173Khz ripple (amplitude of .09K).
08may05: p2030 data feb05 thru apr05 shows aerostat compression with filterbank amps.
11apr05: p2030 pulsar data show compression with new rf amps and wapps are not starting together.
08apr05: totpwr images a2010 drift scans (apr05) show when faa 1350 radar problem occured.
27mar05: 1350 faa radar appears at 1370 to 1390 Mhz.
24mar05: 1430 Mhz birdie in pixel0b of alfa.
28feb,03mar05: faa radar causes increase in power with slow recovery time.
16jan05: beammaps on 3C138 using calibration scan data.
29nov04: calibration scans on B1756+134. Beammaps and Measurement of the zero offset in the rotation angle.
06oct04: faa, punta salinas, and aerostat radars pwr vs time using alfa.
28sep04: a1956 test data. Start of scan ra,dec hdr positions off.
28sep04: compression of alfa receiver from faa radar. Az dependence (64Usecsampling)
24sep04: radar causes power jumps in some alfa beams.
08sep04: 1380 birdie in pixel0b of alfa.
08sep04: 174 khz ripple in pixel4a alfa data.
06sep04: jumps in the total power while doing drift scans.
01aug04: compression from the faa radar.
22apr04: resonance in alfa receiver at 1537.7 Mhz.
22apr04: alfaTsys/Tcal using the old if/lo system.
apr04: Alfa cal values measured in the lab.
HIPASS paper: The HI Parkes all sky survey: southern observations, calibration and robust imaging (pdf file)


Pre Installation (till 21apr04)  (top)

12apr04: computing the wedge position on the floor.
csiro final mount position for alfa (.pdf)
04mar04: testing the alfa rf/if processor with noise sources in the receiver room.
15may03: looking at the alfa IF using lbw horn.
03oct02: Some notes on the upstairs IF power levels that will be needed.
jun02: Field of view measurements using turret offsets and lbn feed.

Do we have to rotate the feed array.
Horn spacing, rotation angle for drift scans, and sampling on the sky.
maximum beam separation for multi beam nyquist sampling with ...german cortes..




 home_~phil